A proton and an alpha particle attempt to penetrate a rectangular potential barrier of height
10 MeV and thickness 10-14m. Both particles have total energies of 5MeV. Use qualitative
arguments to predict which particle has the highest probability of succeeding and evaluate
quantitatively the probability of success of both particles.
1
Expert's answer
2020-06-15T09:52:34-0400

As, both particles have equal energies and "K= \frac {p^2}{2m}" ', thus. the momentum of the alpha particle (since, more massive particle than proton) will be smaller, and the velocity of the proton will be greater. More is the velocity ,thus proton is more likely to get penetrate the barrier.

Now, we know that,


"k_{p}a = \frac{\sqrt{2m_p(V_0 − E)}}{\hbar}a\\ k_{\alpha} a=\frac{\sqrt{2m_{\alpha}(V_0 − E)}}{\hbar}a"

"k pa= \frac {\sqrt (2(938.3MeV/c^ 22)(5 MeV))} {0.6582×10 −15eV⋅s}(10 −14m)=4.9089"

Since, mass of α particle is "m \n_\u03b1\n\u200b\t\n \u22484m_\np\n\u200b"

"k_ \n\u03b1\n\u200b\t\n a=2k_\np\n\u200b\t\n a=9.8179"

Now, we have to calculate the transmission coefficient for both particles to calculate the probability,Hence

"T=(1+ \frac{sinh ^2(ka)}{4 \frac {E }{V _0}(1− V\frac {E }{V _0})})^{-1}"

Thus, substitute value , we get

"T \n_p=2.17869438\u00d710^{\u22124}"

"T _\u03b1=1.1866985\u00d710^{\u22128}"

Thus, corresponding probability is

"0.00217\%\\ 0.000000118\%"

respectively.